OpenStudy (anonymous):

ln(x+1)-ln x=ln 4 (can anyone help me with this)

OpenStudy (anonymous):

if your raise the lns to the e it gets rid of the ln so it comes out to (x+1)-x=4

OpenStudy (anonymous):

ok, you get, x + 1 / x = 4

OpenStudy (anonymous):

can you show me how you solved it

OpenStudy (anonymous):

then you get

OpenStudy (anonymous):

1 + 1/x = 4

OpenStudy (anonymous):

4 = 1/x

OpenStudy (anonymous):

i see you got the answer but can you break it down to me by writting it out so ill understand how to get the answer

OpenStudy (anonymous):

x = 1/4

OpenStudy (anonymous):

i don't have a scanner with me, ok, but here goes

OpenStudy (anonymous):